Using the following information, what is the amount of cost of goods sold?

Purchases $32,000
Merchandise inventory, September 1 5,700
Merchandise inventory, September 30 6,370
Purchase returns and allowances 1,200
Purchase discounts 960
Transportation-in 1,040
Sales 63,000
Sales returns and allowances 910

a.$30,210

b.$28,130

c.$26,900

d.$20,530
Answer #1

Cost of Goods Sold = Beginning inventory + Net Purchases - Ending inventory

Net purchases = 32,000-1200-960+1040 = 30,880

Cost of Goods Sold= 5,700+30,880-6,370

= $30,210 (Answer) Option A